Gerald who is twelve years old is three times as old as his sister Jane how old will Gerald be when he is twice as old as Jane.. 13;16;17;19 or 22
Answer by stanbon(75887)

Gerald who is twelve years old is three times as old as his sister Jane how old will Gerald be when he is twice as old as Jane.. 13;16;17;19 or 22
--------
G = 12 ; J = 4
----
Equation:
12+x = 2(4+x)
12+x = 8 + 2x
x = 4 yrs
-------
G+4yrs = 16 yrs old
J+4yrs = 8 yrs old
----------------
Cheers,
Stan H.
----------